40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1720     Entry time: Thu Sep 22 11:14:36 2016
Author: awade 
Type: Summary 
Category: TempCtrl 
Subject: Thermal controls: cavity shields and vacuum tank (draft schematic)  

I've put together a draft schematic of the basic layout of the shields and tank as well as the present state of connections of the dual ref cavity setup (attached in both pdf and omnigraffle formate).  There are presently no active controls on either the individual cavity shields nor the whole vacuum can. We are relying on passive stability. 

The tank heaters were hooked up to a large current Hewlett Packard 6267B DC power supply.  The unit is no longer producing voltage or current, either a fuse is blown or its kaput; I have removed the unit from the rack. There are a few more 6267Bs around the lab and all have similar problems. I have stacked them all in the corner for assessment and possible disposal in the next big lab cleanup.

Resistive temperature sensors on the two shields are both broken while resistive heating wires on both shields are working. However, according to Tara the length of wire is different because they didn't know at the time if they would be voltage or current limited in their actuation: to mitigate for both cases they chose for the south path to have 85.6Ω resistance and the north 156.8Ω. When we install the new shields we will bring these to be identical. 

I'm not really sure about the existence of a voltage-to-current drivers used in the past.  Its evident that there was one for the previous implementation of PID control of the vacuum can, but I haven't seen any boxes/circuitry in the lab. We need a low noise high current driver for our new implementation. Aidan has linked to the TCS drivers in PSL:1711. Our design requirements will come down to the characteristics of the plant and how it filters the LF noise.

 

Some information we don't have yet:

  • The mass of the tank;
  • the type of temperature sensors used on the current shields;
  • a model of radiative transfer between shields and the tank.  This is significant complicated by the fact that the Al frame and mechanical isolation stages are actually acting as a shield as well (they go all the way around);
  • the schematic of the AD590 -> voltage adaptor circuitry.  The LIGO DCC number appears to be for a VCO servo even though it is labeled "LIGO Temperature Sensor Interface board" a search of the DCC or elog doesn't seem to turn up anything;
  • What current drivers were used in the past to drive the vacuum can heaters
  • Modeling of full impact of end caps? 
  • modeling of improvement from moving to 1" windows on vac can?

 

 

Note: the omnigraffle file has multiple canvases with different info. 

Attachment 1: ThermalControlsDiagram_v8.graffle.zip  124 kB
Attachment 2: ThermalControlsDiagram_v8.pdf  155 kB  | Hide | Hide all
ThermalControlsDiagram_v8.pdf ThermalControlsDiagram_v8.pdf ThermalControlsDiagram_v8.pdf
ELOG V3.1.3-